OpenStudy (oliviagggg):

What is the next term in the sequence 1/8,1/4,1/2,...

OpenStudy (oliviagggg):

1 1 1 - - - ..... 8 4 2

OpenStudy (tkhunny):

There are infinitely many correct responses to this question as it is presented. Is it a GEOMETRIC sequence? That would narrow it down.

OpenStudy (oliviagggg):

it is geometric

OpenStudy (anonymous):

The answer is 1

OpenStudy (anonymous):

It is multiplying by two. 1/8 ×2 = 1/4 1/4 ×2 = 1/2 1/2 × 2 = 1

OpenStudy (oliviagggg):

i thought it was 1 too but i wasnt sure Thank you

OpenStudy (tkhunny):

You must DECLARE that it is GEOMETRIC. It may not be assumed.

OpenStudy (oliviagggg):

it said it was on my question i just forgot to put it i guess

OpenStudy (oliviagggg):

Thanks :)